# TropNNC: Structured Neural Network Compression Using Tropical Geometry

### v2

@@ 7.5 Non-uniform Pruning
Original - 11.69M 3.64G 70.10
CUP 0.5 11.66M 3.58G 66.60
TropNNC (v2) 1.1 11.64M 3.44G 68.40
CUP 0.6 11.49M 3.38G 53.20
TropNNC (v2) 1.2 11.27M 3.07G 57.90
CUP 0.65 11.17M 3.15G 29.20
TropNNC (v2) 1.25 10.71M 2.75G 41.20
TropNNC (v1) 0.021 9.87M 3.46G 58.10
8 Conclusions and Future work
We improved upon the theoretical results of Misiakos et al. (2022) by using the (usual) Hausdorff distance instead of its discrete counterpart, to obtain a tighter bound for the approximation of tropical polynomials. Leveraging this enhancement we refined the choice of compressed node’s weights that correspond to the K-means clusters from (Misiakos et al., 2022). Our novel algorithm, TropNNC, applied to linear and convolutional layers of deep neural networks outperforms Neural Path K-means of Misiakos et al. (2022), and manages to match or even surpass the performance of ThiNet, particularly for the compression of linear layers. In non-uniform pruning it performs better than CUP, with significant improvement in the case of the VGG network architecture. These achievements underscore the potential of tropical geometry in the realm of neural network compression. One promising direction for further investigation could involve optimizing the approximation techniques by considering the upper hulls of the network’s tropical polynomials.
Appendix A Proofs of Propositions and Lemmas
A.1 Proof of Lemma 2
@@
Proof  Since $(\mathbf{a}^{T},b)$ lies below the upper envelope of $P$, there exists a point
@@
$=\sum_{i=1}^{k}\lambda_{i}\langle\mathbf{v}_{i},(\mathbf{x},1)\rangle\leq\max_{i=1,\ldots,k}\langle\mathbf{v}_{i},(\mathbf{x},1)\rangle\leq p(\mathbf{x}).$
If $(\mathbf{a}_{i}^{T},b_{i})$ lies strictly below the upper envelope of $P$, then $b^{\prime}>b$, and the inequality is strict.
A.2 Proof of Lemma 8
Proof  We will prove that
$\sup_{\mathbf{u}\in P}\operatorname{dist}(\mathbf{u},\tilde{P})=\max_{\mathbf{u}\in V_{P}}\operatorname{dist}(\mathbf{u},\tilde{P}),$
@@ A.2 Proof of Lemma 8
which concludes the proof.
A.3 Proof of Proposition 10 and Theorem 11
Proof  Consider a point $\mathbf{x}\in B$ and assume that $p(\mathbf{x})=\mathbf{a}^{T}\mathbf{x}+b$ and $\tilde{p}(\mathbf{x})=\mathbf{c}^{T}\mathbf{x}+d$. Take an arbitrary $(\mathbf{u}^{T},v)\in\tilde{P}$. This point lies below the upper envelope of $\tilde{P}$. Thus, by Lemma 2, we have that $\tilde{p}(\mathbf{x})\geq\mathbf{u}^{T}\mathbf{x}+v$. Choose $(\mathbf{u}^{T},v)$ to be the closest point to $(\mathbf{a}^{T},b)$. Then,
@@ A.3 Proof of Proposition 10 and Theorem 11
$-\max_{(\mathbf{c}^{T},d)\in V_{\tilde{P}}}\operatorname{dist}((\mathbf{c}^{T},d),P)\cdot\rho\leq p(\mathbf{x})-\tilde{p}(\mathbf{x})\leq\max_{(\mathbf{a}^{T},b)\in V_{P}}\operatorname{dist}((\mathbf{a}^{T},b),\tilde{P})\cdot\rho,$
@@
$\frac{1}{\rho}\max_{\mathbf{x}\in B}\lvert p(\mathbf{x})-\tilde{p}(\mathbf{x})\rvert\leq H(P,\tilde{P}),$
which concludes the proof of Proposition 10.
Finally, notice that we may write
$\|\mathbf{v}(\mathbf{x})-\tilde{\mathbf{v}}(\mathbf{x})\|_{1}$ $=\sum_{j=1}^{m}\lvert\mathbf{v}_{j}(\mathbf{x})-\tilde{\mathbf{v}}_{j}(\mathbf{x})\rvert$
$=\sum_{j=1}^{m}\lvert(p_{j}(\mathbf{x})-q_{j}(\mathbf{x}))-(\tilde{p}_{j}(\mathbf{x})-\tilde{q}_{j}(\mathbf{x}))\rvert$
$=\sum_{j=1}^{m}\lvert(p_{j}(\mathbf{x})-\tilde{p}_{j}(\mathbf{x}))-(q_{j}(\mathbf{x})-\tilde{q}_{j}(\mathbf{x}))\rvert$
$\leq\sum_{j=1}^{m}\left(\lvert p_{j}(\mathbf{x})-\tilde{p}_{j}(\mathbf{x})\rvert+\lvert q_{j}(\mathbf{x})-\tilde{q}_{j}(\mathbf{x})\rvert\right).$
By Proposition 10 we derive
$\frac{1}{\rho}\max_{\mathbf{x}\in B}\|\mathbf{v}(\mathbf{x})-\tilde{\mathbf{v}}(\mathbf{x})\|_{1}\leq\sum_{j=1}^{m}\left(H(P_{j},\tilde{P}_{j})+H(Q_{j},\tilde{Q}_{j})\right).$
A.4 Proof of Proposition 12
Proof  For the output function, it holds that
@@ A.4 Proof of Proposition 12
$\frac{1}{\rho}\left|v(\mathbf{x})-\tilde{v}(\mathbf{x})\right|\leq H(P,\tilde{P})+H(Q,\tilde{Q}).$
Thus, it suffices to get a bound on $H(P,\tilde{P})$ and $H(Q,\tilde{Q})$.
Let $I_{+}\subseteq[n]$ and $I_{-}\subseteq[n]$ be the sets of positive and negative generators, respectively. First, we deal with $H(P,\tilde{P})$. Notice that $\forall\mathbf{v}\in V_{\tilde{P}},\mathbf{v}$ is the sum of generators of some clusters of $P$. Thus $\forall\mathbf{v}\in V_{\tilde{P}},\mathbf{v}$ is a vertex of $P$. Hence,
@@
Thus, for every cluster $k\in C_{+}$, the point $\tilde{x}_{k}\tilde{c}_{k}(\tilde{\mathbf{a}}_{k}^{T},\tilde{b}_{k})$ lies inside the segment $[\mathbf{0},\tilde{c}_{k}(\tilde{\mathbf{a}}_{k}^{T},\tilde{b}_{k})]$ and thus belongs to $\tilde{P}$.
@@
$\operatorname{dist}(\mathbf{u},\tilde{P})$ $\leq\left\|\sum_{i\in I_{+}^{\prime}}c_{i}(\mathbf{a}_{i}^{T},b_{i})-\sum_{k\in C_{+}}\tilde{x}_{k}\tilde{c}_{k}(\tilde{\mathbf{a}}_{k}^{T},\tilde{b}_{k})\right\|$
@@
$H(P,\tilde{P})$ $=\max\left\{\max_{\mathbf{u}\in V_{P}}\operatorname{dist}(\mathbf{u},\tilde{P}),\max_{\mathbf{v}\in V_{\tilde{P}}}\operatorname{dist}(P,\mathbf{v})\right\}$
@@
Similarly, for $H(Q,\tilde{Q})$, we have
@@
which concludes the proof of Proposition 12.
A.5 Proof of Corollary 13
Proof  The bound of Misiakos et al. (2022) is the following:
@@ A.5 Proof of Corollary 13
which concludes the proof.
A.6 Proof of Proposition 14
@@
Definition 15 (Null neuron/generator)
@@
Proof  Assume the algorithm iterative scheme has reached a stationary point (otherwise, assume the last step is an output weight update and the proof works fine). First, we focus on a single output, say $j$-th output. We will bound $H(P_{j},\tilde{P}_{j}),H(Q_{j},\tilde{Q}_{j})$ for all $j\in[m]$.
@@
Thus, for every cluster $k\in C_{j+}$, the point $\tilde{x}_{jk}\tilde{c}_{jk}(\tilde{\mathbf{a}}_{k}^{T},\tilde{b}_{k})$ lies inside the segment $[\mathbf{0},\tilde{c}_{jk}(\tilde{\mathbf{a}}_{k}^{T},\tilde{b}_{k})$], and thus belongs to $\tilde{P_{j}}$.

### v3

@@ 6.3 Non-uniform Pruning
Original - 14.7M 0.63G 93.64
TropNNC 0.02 2.75M 0.39G 86.2
TropNNC (3 iters) 0.021 2.52M 0.38G 91.2
TropNNC 0.025 1.80M 0.33G 34.89
TropNNC (3 iters) 0.0255 1.76M 0.33G 69.95
7 Conclusion
We proposed TropNNC, a tropical geometrical method for structured, data-free pruning of linear and convolutional layers of ReLU activated neural networks. Based on clustering similar neurons/channels, it selects better cluster representatives than previous work by using tropical geometry and optimization, with provable compression guarantees. TropNNC significantly outperforms prior work on tropical geometric pruning and manages to match or even surpass the performance of the data-driven ThiNet. In non-uniform pruning it outperforms CUP, with significant improvement in the case of the VGG architecture. Our findings highlight the potential of tropical geometry in the realm of neural network compression.
Acknowledgments
The research project was supported by the Hellenic Foundation for Research and Innovation (H.F.R.I.) under the “2nd Call for H.F.R.I. Research Projects to support Faculty Members & Researchers” (Project Number:2656, Acronym: TROGEMAL).
Appendix
In this appendix, we first go into more detail regarding the heuristic improvements of our algorithm. Afterwards, we introduce the background of tropical algebra which was omitted from the main text. Finally, we provide the proofs for the theorems and propositions stated in the main text.
Before we proceed with the technical content of this appendix, we first provide an ethical statement for this work and a disclosure of the limitations of our method.
Ethical statement
This work does not introduce any new ethical concerns beyond those already present in prior research on network pruning and model compression. As with existing pruning methods, our approach aims to reduce model size and computational cost, and does not involve the use of sensitive data, human subjects, or deployment in high-risk applications.
Limitations
A key limitation of our method is that it relies on the presence of consecutive convolutional or linear layers – that is, the existence of intermediate hidden layers. In addition, it only holds for networks with ReLU activations.
Heuristic Improvements
In the experiments comparing TropNNC and CUP, TropNNC was enhanced with 2 heuristic improvements. These improvements are not rigorously backed by theory; however, we try to provide intuitive explanations as to why they work.
Heuristic 1.
First, let us focus on the single-output case, where the problem of network compression reduces to the problem of zonotope approximation. In this setting, Algorithm 1 suggests clustering similar generators based on their $L^{2}$ distances and replacing each cluster with the sum of its elements. The intuition is as follows: if the generators within a cluster are exactly equal, then through their Minkowski sum, these generators become equivalent to a single generator given by their sum. We can take this one step further: if we have generators that are parallel to each other, their Minkowksi sum reduces to a single equivalent generator without introducing any error. This suggest that generator similarity is determined by direction (i.e. cosine similarity) and not $L^{2}$ distance. Hence, an improvement to the single-output algorithm is to cluster the generators based on cosine similarity. In practice, this can be achieved by applying K-Means clustering to normalized generators, while still using the original generators in the summation step.
A similar heuristic can also be applied to the multi-output algorithm. By normalizing the input weights of the clustering vectors (i.e. the vectors used for clustering similar neurons/channels, not the weights used for calculating the representative of each cluster) we obtain an improvement to the algorithm.
Heuristic 2.
The second heuristic is motivated by Theorem 9, which states that two tropical polynomials are functionally identical if the upper hulls of their extended Newton polytopes coincide (i.e. only the vertices of the upper hull contribute to non-redundant terms). Indeed, this theorem also serves as motivation for Theorem 2, which shows that approximate equality of the extended Newton polytopes implies approximate functional equivalence of the corresponding tropical polynomials. Although we do not formally prove it, this theorem can be further refined by bounding the error of two tropical polynomials by the Hausdorff distance of the upper hulls of their extended Newton polytopes:
$)\frac{1}{\rho}\max_{\mathbf{x}\in B}|p(\mathbf{x})-\tilde{p}(\mathbf{x})|\leq H(UF(P),UF(\tilde{P}$))
Based on this refined result, it becomes clear that to compress networks, it suffices to approximate the upper hulls of the zonotopes of the network, rather than the zonotopes themselves.
We can take this idea one step further. Consider the upper hull of an extended Newton polytope, and project it by eliminating the bias term. Then, the points on the convex hull of the resulting projection (which coincides with the convex hull of the (non-extended) Newton polytope) correspond to the slopes of the outer linear regions of the polynomials, i.e. the regions which extend to infinity. We can view these points as being the "outer" points of the upper hull. When approximating a tropical polynomial, these unbounded regions are particularly important, because they have the potential to make the sup-norm of the approximation error infinite (something which is not captured by Theorem 2 because we divide by $\rho$ and take the $L^{1}$-norm error on a bounded set).
Based on this insight, our goal is to design an algorithm that prioritizes the approximation of the upper hulls of the zonotopes – particularly the vertices corresponding to the unbounded regions of the tropical polynomials. (i.e. the "outer" vertices of the upper hulls).
Let us first see how vanilla TropNNC behaves on a simple example. Consider the tropical polynomial $v(x)=\max(0,-x+5)+\max(0,x+5)+\max(0,x)$. Such a polynomial and its corresponding zonotope are depicted in Figure 6(a). Notice that the upper hull has 4 points, each corresponding to a linear region of the polynomial. It also has 2 "outer" points, which correspond to the 2 outer regions of the polynomial extending to infinity.
We reduce the terms from 3 to $K=2$. First, TropNNC performs clustering on the generators of the zonotope, forming the clusters $\{(-1,5),(1,5)\}$ and $\{(1,0)\}$. It then replaces the first cluster with the representative $(0,10)$ and the second cluster with the representative $(1,0)$. The reduced tropical polynomial is $\tilde{v}(x)=\max(0,10)+\max(0,x)=10+\max(0,x)$. The compression procedure and the corresponding reduced function is depicted in Figure 6(b).
It is evident that this approximation is not ideal: while the reduced polynomial accurately captures the two central regions, it fails to preserve the unbounded regions. This causes the approximation error becomes unbounded.
Since we have $K=2$, the best possible approximation using only 2 terms is $v^{\prime}(x)=\max(0,-x+5)+\max(0,2x+5)$. This approximation captures both the important outer regions with zero error, and also one of the central regions. The question however remains: how could we have arrived at such an approximation? To answer this, notice that $v^{\prime}(x)=\max(0,-x+5)+\max(0,(x+5)+(x))$. Thus, if TropNNC had initially formed the clusters $\{(-1,5)\}$ and $\{(1,5),(1,0)\}$, then it would have been able to find the better approximation. Now, the modification we have to make becomes clear: instead of forming the clusters based on the generators themselves, we can form the clusters based on the reduced generators with the bias suppressed (and, of course, for the summation step the bias remains). The modified compression procedure is depicted in Figure 6(c). The new approximation clearly provides a closer fit to the upper hull and captures the “outer” points exactly, resulting in zero error in the outer unbounded regions.
To conclude, another heuristic improvement comes from dropping the bias from the clustering vectors, seeing as this helps the algorithm achieve better upper hull approximations and prioritize the important "outer" vertices of the upper hull.
(a) Example tropical polynomial $v(x)=\max(0,-x+5)+\max(0,x+5)+\max(0,x)$ and corresponding zonotope.
(b) Approximation using vanilla TropNNC.
(c) Approximation using Heuristic 2.
Figure 6: Example of reduction of tropical polynomial.
Background on Tropical Algebra
We begin with the definitions of tropical polynomials, tropical rational functions, Newton polytopes, and zonotopes.
Tropical Polynomials and Rational Functions.
where $n]:=\{1,...,n\}$. Here, $n$ represents the rank of the tropical polynomial. Each monomial term $\{\mathbf{a}_{i}^{T}\mathbf{x}+b_{i}\}$ of the polynomial has an exponent or slope $\mathbf{a}_{i}\in\mathbb{R}^{d}$ and a coefficient or bias $b_{i}\in\mathbb{R}$. Each monomial term corresponds to a plane in $\mathbb{R}^{d+1}$. Consequently, tropical polynomials are piecewise linear convex functions. Specifically, every tropical polynomial is a continuous piecewise linear convex function, and every continuous piecewise linear convex function can be expressed (though not uniquely) as a tropical polynomial (Maclagan and Sturmfels, 2021). The set of tropical polynomials in $\mathbf{x}$ defines the semiring $\mathbb{R}_{\mathrm{max}}[\mathbf{x}]$. Figures 7(a), 7(b) illustrate examples of tropical polynomials in one and multiple variables, respectively.
Tropical rational functions are defined as the difference of two tropical polynomials $p$ and $q$:
$r(\mathbf{x})=p(\mathbf{x})-q(\mathbf{x})$
Tropical rational functions correspond to general piecewise linear functions. Specifically, every tropical rational function is a continuous piecewise linear function, and every continuous piecewise linear function can be expressed (though not uniquely) as a tropical rational function. Figures 7(c), 7(d) provide examples of tropical rational functions.
Newton Polytopes.
For a tropical polynomial $f$ as defined in (4), we define its Newton polytope as the convex hull of the slopes $\mathbf{a}_{i}$ of $f$
$\mathrm{Newt}(f):=\mathrm{conv}\left\{\mathbf{a}_{i},i\in[n]\right\}.$
The following proposition allows us to calculate the (extended) Newton polytope of expressions of tropical polynomials, and it was used in the main text to deduce that the extended Newton polytopes of the outputs of the network are zonotopes.
Proposition 7 (Zhang et al., 2018).
$\mathrm{ENewt}(f+g)$ $=\mathrm{ENewt}(f)\oplus\mathrm{ENewt}(g).$
The Minkowski sum $\oplus$ of two polytopes (or more generally subsets of $\mathbb{R}^{d}$) $P$ and $Q$ is defined as:
$P\oplus Q:=\{p+q\mid p\in P,q\in Q\}.$
The upper envelope or upper hull $UF(P)$ of an extended Newton polytope is defined as the set of all points $(\mathbf{a}^{T},b)$ of the polytope $P$ that are not "shadowed" by any other part of the polytope when viewed from above (last dimension). This means that there is no $b^{\prime}>b$ such that $(\mathbf{a}^{T},b^{\prime})$ belongs to $P$. We have the following useful lemma.
Lemma 8.
Proof.
@@ Acknowledgments
Since $(\mathbf{a}^{T},b)$ lies below the upper envelope of $P$, there exists a point
@@
$\sum_{i=1}^{k}\lambda_{i}\langle\mathbf{v}_{i},(\mathbf{x},1)\rangle\leq\max_{i=1,\ldots,k}\langle\mathbf{v}_{i},(\mathbf{x},1)\rangle\leq p(\mathbf{x}).$
If $(\mathbf{a}_{i}^{T},b_{i})$ lies strictly below the upper envelope of $P$, then $b^{\prime}>b$, and the inequality is strict. ∎
The extended Newton polytope provides a geometrical interpretation for studying tropical polynomials. For instance, the following theorem holds.
Theorem 9.
$f=g\Leftrightarrow UF(\mathrm{ENewt}(f))=UF(\mathrm{ENewt}(g))$
Example 2.
Consider the polynomials corresponding to Figure 8.
$f(x,y)=\max\{0,-y+1,y+1\}$
$g(x,y)=\max\{x+1,-x+1\}.$
We have that
$(f\vee g)(x,y)=\max\{0,x+1,y+1,-y+1,-x+1\}$
$(f+g)(x,y)=\max\{x+1,-x+1,x-y+2,$
xy+2,x+y+2,x+y+2}.-x-y+2,x+y+2,-x+y+2\}.
$(f\vee g)(x,y)=\max\{x+1,y+1,-y+1,-x+1\},$
(a) $f$
(b) $g$
(c) $f\vee g$
(d) $f+g$
Zonotopes.
$Z:=\bigoplus_{i\in[n]}g_{i}$
The line segments $g_{i}$ are referred to as the zonotope’s generators.
Alternatively, a zonotope can be expressed equivalently by a set of vectors $\mathbf{v}_{1},\ldots,\mathbf{v}_{n}\in\mathbb{R}^{d}$ and a starting point $\mathbf{s}\in\mathbb{R}^{d}$. By taking the generators to be the segments $\mathbf{0},\mathbf{v}_{1}],\ldots,[\mathbf{0},\mathbf{v}_{n}]$ and translating the first segment by $\mathbf{s}$, we obtain the equivalent form:
In this context, the vectors $\mathbf{v}_{i}$ are sometimes referred to as the zonotope’s generators, meaning the segments $\mathbf{0},\mathbf{v}_{i}]$. When the starting point $\mathbf{s}$ is not mentioned, it is assumed to be the origin $\mathbf{0}$.
$\mathbf{u}=\mathbf{s}+\sum_{i\in I}\mathbf{v}_{i},$
where $I\subseteq[n]$.
Tropical polynomial approximation based on Hausdorff distance
Before we proceed with the proof of our refined Theorem 2, we first provide an auxiliary lemma, stated implicitly in the main text, which is necessary to see how our bound is indeed tighter than the bound of Misiakos et al. (2022). This lemma is also used for the proofs of Theorem 2, and Propositions 4 and 6.
Lemma 10.
Due to the convexity and compactness of polytopes, we have that
Proof.
We will prove that
[…]
@@
which concludes the proof. ∎
We continue with the proof of Theorem 2.
Proof (Theorem 2).
Consider a point $\mathbf{x}\in B$ and assume that $p(\mathbf{x})=\mathbf{a}^{T}\mathbf{x}+b$ and $\tilde{p}(\mathbf{x})=\mathbf{c}^{T}\mathbf{x}+d$. Take an arbitrary $(\mathbf{u}^{T},v)\in\tilde{P}$. This point lies below the upper envelope of $\tilde{P}$. Thus, by Lemma 8, we have that $\tilde{p}(\mathbf{x})\geq\mathbf{u}^{T}\mathbf{x}+v$. Choose $(\mathbf{u}^{T},v)$ to be the closest point to $(\mathbf{a}^{T},b)$. Then,
@@
$\max_{(\mathbf{c}^{T},d)\in V_{\tilde{P}}}\operatorname{dist}((\mathbf{c}^{T},d),P)\cdot\rho\leq p(\mathbf{x})-\tilde{p}(\mathbf{x})\leq\max_{(\mathbf{a}^{T},b)\in V_{P}}\operatorname{dist}((\mathbf{a}^{T},b),\tilde{P})\cdot\rho,$
@@
[…]
Proofs of Proposition 4 and Corollary 5
We continue with the proofs of Proposition 4 and Corollary 5.
Proof (Prop. 4).
For the output function, it holds that
@@
[…]
Thus, it suffices to get a bound on H(P,P~)H($\tilde{P}$) and H(Q,Q~)H($\tilde{Q}$).
Let $I_{+}\subseteq[n]$ and $I_{-}\subseteq[n]$ be the sets of positive and negative generators, respectively. First, we deal with H(P,P~)H($\tilde{P}$). Notice that $\forall\mathbf{v}\in V_{\tilde{P}},\mathbf{v}$ is the sum of generators of some clusters of $P$. Thus $\forall\mathbf{v}\in V_{\tilde{P}},\mathbf{v}$ is a vertex of $P$. Hence,
@@
Thus, for every cluster $k\in C_{+}$, the point $\tilde{x}_{k}\tilde{c}_{k}(\tilde{\mathbf{a}}_{k}^{T},\tilde{b}_{k})$ lies inside the segment $\mathbf{0},\tilde{c}_{k}(\tilde{\mathbf{a}}_{k}^{T},\tilde{b}_{k})]$ and thus belongs to $\tilde{P}$.
@@
[…]
@@
[…]
@@
Similarly, for H(Q,Q~)H($\tilde{Q}$), we have
@@
which concludes the proof of Proposition 4. ∎
Proof (Corollary 5).
The bound of Misiakos et al. (2022) is the following:
@@
which concludes the proof. ∎
Proof of Proposition 6
@@
Definition 11 (Null neuron/generator).
@@
Proof (Prop. 6).
Assume the algorithm iterative scheme has reached a stationary point (otherwise, assume the last step is an output weight update and the proof works fine). First, we focus on a single output, say $j$-th output. We will bound $H(P_{j},\tilde{P}_{j}),H(Q_{j},\tilde{Q}_{j})$ for all $j\in[m]$.
@@
Thus, for every cluster $k\in C_{j+}$, the point $\tilde{x}_{jk}\tilde{c}_{jk}(\tilde{\mathbf{a}}_{k}^{T},\tilde{b}_{k})$ lies inside the segment $\mathbf{0},\tilde{c}_{jk}(\tilde{\mathbf{a}}_{k}^{T},\tilde{b}_{k})$], and thus belongs to $\tilde{P_{j}}$.